find the distance between p1 and p2
p1=(5,-2)
p2=(4,5)

Answer by stanbon(75887)

find the distance between p1 and p2
p1=(5,-2)
p2=(4,5)
------
d = sqrt[(5--2)^2 + (5-4)^2] = sqrt[49+1] = 5*sqrt(2)
===================
Cheers,
Stan H.
=============